# The Resolution of the Gauge Problem of Cosmology provides Insight into the Formation of the First Structures in the Universe

### v2

@@ XII.2 Era after Decoupling of Matter and Radiation
In these expressions, the symbol $k_{\text{B}}$ represents Boltzmann’s constant, the quantity $m$ denotes the proton rest mass, and the temperature of the matter is indicated by the symbol $T$. The rest mass energy density is $nmc^{2}$ and the kinetic energy density is $\tfrac{3}{2}nk_{\text{B}}T$. The elimination of the temperature, $T$, from the equations of state (88), yields the equation of state for the pressure [see Ref. [31], Eq. (2.10.27)]:
@@
The partial derivatives are determined by the utilization of the expressions (13b) and are given by
@@
$\quad\beta\approx\dfrac{v_{\text{s}}}{c}=\sqrt{\dfrac{5}{3}\dfrac{k_{\text{B}}T_{\scriptscriptstyle(0)}}{mc^{2}}}.$ (92)
In this expression, $v_{\text{s}}$ represents the adiabatic speed of sound. From $\beta^{2}\approx\tfrac{5}{3}w$, it is found that Eq. (44) results in $\dot{w}\approx-2Hw$. Therefore, with $H:=\dot{a}/a$, it can be concluded that $w\propto a^{-2}$. Given that $w$ is proportional to $T_{{\scriptscriptstyle(0)}}$, the well-known result (see Ref. [31], Eq. (15.5.16) with $\gamma=\tfrac{5}{3}$, where ${\gamma:=c_{\text{p}}/c_{\text{v}}}$ is the adiabatic index) is obtained:
@@
In light of the relatively minor values of both $w$ and $\beta^{2}$, it is now possible to justify the equations of state (88). From the proportionality (93) and the fact that $n_{{\scriptscriptstyle(0)}}\propto a^{-3}$, as can be seen from Eq. (71c), it follows from (88) that ${p_{{\scriptscriptstyle(0)}}\propto a^{-5}}$. Given that pressure decayed rapidly and was already negligible in comparison to the energy density at the time of decoupling, it can be concluded that the unperturbed pressure had a minimal impact on structure formation. It is therefore reasonable to neglect the quantities $w$ and $\beta^{2}$ with respect to the constants of order one that appear in the momentum conservation laws (45). In doing so, one finds
The second equation is derived by lowering the index $i$ in the first equation using the background metric $g_{{\scriptscriptstyle(0)}ij}$ and the relation $\dot{g}_{{\scriptscriptstyle(0)}ij}=2Hg_{{\scriptscriptstyle(0)}ij}$. Accordingly, the evolution of density fluctuations is highly dependent upon the presence of pressure gradients, represented by $p_{{\scriptscriptstyle(1)}|j}$. In the event that the pressure gradients are identically zero, Eqs. (94) reduce to the form of Eqs. (49) and (50) in the nonrelativistic limit. This implies that the physical component of the fluid velocity is zero, and consequently, density perturbations cease to evolve. Given that structure formation in the universe was predominantly determined by pressure gradients, it is necessary to consider the kinetic energy density and pressure in addition to the energy density. This is the reason why the particle number density must be taken into account in the equations of state. In conclusion, it can be inferred from the equations (94) that the gauge modes (28) are zero. This result demonstrates that the relativistic gauge transformations (1), where $\xi^{\mu}$ is given by (6), are equivalent to the Newtonian gauge transformations (53). This conclusion is consistent with expectations, given that the equations of state (88) are nonrelativistic.
The subsequent step is to derive the evolution equations for density perturbations. The proportionality (93) implies with (92) that $\dot{\beta}/\beta=-H$. The system of equations (60) with coefficients (72) can now be rewritten in the following form:
@@
In this calculation, the approximation is made that $w$ and $\beta^{2}$ are negligible with respect to the constants of order one.
@@
$\delta_{n}-\delta_{\varepsilon}\propto a^{-2}.$ (96)
In order to relate this proportionality to thermodynamic quantities, it is necessary to express the perturbed equation of state for the energy density in terms of density fluctuations. It can be deduced from $\varepsilon=\varepsilon(n,T)$ that
@@
$\varepsilon_{{\scriptscriptstyle(1)}}^{{\textnormal{phys}}}=\left(\dfrac{\partial\varepsilon}{\partial n}\right)_{T}n_{\scriptscriptstyle(1)}^{{\textnormal{phys}}}+\left(\dfrac{\partial\varepsilon}{\partial T}\right)_{n}T_{\scriptscriptstyle(1)}^{{\textnormal{phys}}}.$ (98)
This result is based on the use of expressions (4a) and (68). By employing the expression for $\varepsilon$ in (88) to eliminate the partial derivatives, one obtains the perturbed equation of state:
@@
By dividing expression (99) by $\varepsilon_{\scriptscriptstyle(0)}$, and using the exact value of the ratio $w:=p_{{\scriptscriptstyle(0)}}/\varepsilon_{{\scriptscriptstyle(0)}}$, the perturbed equation of state for the energy density expressed in fluctuations (59) is obtained:
@@
In this context, the quantity $\delta_{T}$ is defined by $\delta_{T}:=T^{\textnormal{phys}}_{\scriptscriptstyle(1)}/T_{\scriptscriptstyle(0)}$. The solution to Eq. (95b) is obtained by combining (96) and (100) and using the approximate value of $w$, (91), which implies that $w\propto a^{-2}$. The result is as follows:
@@
The quantity $\delta_{T}(t_{\text{dec}},\bm{x})$ represents the temperature fluctuation of matter at the time $t_{\text{dec}}$, which marks the decoupling of matter from radiation. Since $w\ll 1$, (101) is in alignment with (87).
According to (67), (91) and (101) the entropy per particle is as follows:
@@
Therefore, it can be concluded that the quantity $\delta_{T}$ is random and that density fluctuations are nonadiabatic if $\delta_{T}$ is nonzero.
Upon substituting (90), (92), (101), and the approximation ${w\ll 1}$, the local pressure fluctuation, defined by $\delta_{p}:=p^{\textnormal{phys}}_{\scriptscriptstyle(1)}/p_{\scriptscriptstyle(0)}$, can be calculated from expression (70). One obtains
@@
In this context, the terms $\tfrac{5}{3}\delta_{\varepsilon}(t,\bm{x})$ and $\delta_{T}(t_{\text{dec}},\bm{x})$ represent the adiabatic and nonadiabatic pressure fluctuations, respectively.
The solution to Eq. (95b) has been determined. This allows for the reformulation of the second-order equation (95a) to facilitate analysis of the evolution of density fluctuations. To that end, the solutions to the background equations (71) are required. Given that the pressure with respect to the rest mass energy density can be disregarded ($w\ll 1$), the following solutions are obtained:
@@
The dimensionless time, denoted by the symbol $\tau$, is defined as follows: $\tau:=t/t_{\text{dec}}\geq 1$. This definition implies that
@@
By employing Eqs. (71a) and (101), as well as the expressions (92), (104) (105), the Helmholtz equation $\nabla^{2}\delta=-|\bm{q}|^{2}\delta$, and $w\propto a^{-2}$, it can be demonstrated that Eq. (95a) can be expressed in the following manner:
$\delta_{\varepsilon}^{\prime\prime}+\dfrac{2}{\tau}\delta_{\varepsilon}^{\prime}+\left(\dfrac{4}{9}\dfrac{\mu_{\text{m}}^{2}}{\tau^{8/3}}-\dfrac{10}{9\tau^{2}}\right)\delta_{\varepsilon}=-\dfrac{4}{15}\dfrac{\mu^{2}_{\text{m}}}{\tau^{8/3}}\delta_{T}(t_{\text{dec}},\bm{q}).$ (106)
In this equation, the prime denotes differentiation with respect to the variable $\tau$. The parameter $\mu_{\text{m}}$ is defined as follows:
@@
In this expression, $\lambda_{\text{dec}}:=\lambda a(t_{\text{dec}})$ represents the physical scale of a fluctuation at $t=t_{\text{dec}}$. The magnitude of the fluctuation is given by $\lvert\bm{q}\rvert=2\pi/\lambda_{\text{dec}}$. Furthermore, one has $v_{\text{s}}(t_{\text{dec}})/c\approx\beta(t_{\text{dec}})$, as stated in expression (92).
Finally, the expression (103) enables the reformulation of Eq. (106) in a form suitable for the study of the evolution of density fluctuations:
@@
In the limit of large scales, that is, $\lambda_{\text{dec}}\rightarrow\infty$, which leads to $\mu_{\text{m}}\rightarrow 0$, the solution of Eq. (108) is
@@
Therefore, pressure fluctuations $\delta_{p}$ played a negligible role in the evolution of the large-scale fluctuations. In other words, these fluctuations were almost adiabatic and evolved solely under the influence of gravitation and expansion.
A comparison of the solutions (109) and (157a) reveals that the solution ${\propto\tau^{2/3}}$ is a consequence of the homogeneous part of Eq. (156a). The solution $\propto\tau^{-5/3}$ is the particular solution to the inhomogeneous equation. This solution can only be obtained if the covariant divergence $\vartheta_{{\scriptscriptstyle(1)}}$ is taken into account, as demonstrated in Appendix C.2. In contrast to the standard equation (150), the quantity $\vartheta_{{\scriptscriptstyle(1)}}$ is included in Eq. (95a). The presence of the gauge constant $C$, precludes the derivation of solution (109) from the system (156).
XII.2.1 Cosmological Quantities
To investigate the growth of density fluctuations, it is first necessary to express the parameter $\mu_{\text{m}}$ (107) in observable quantities.
@@ XII.2.1 Cosmological Quantities
where $a(t_{\text{p}})$ is the present value of the scale factor. In the case of a flat flrw universe, it is permissible to take $a(t_{\text{p}})=1$. Using the background solutions (104), it is possible to express the relevant parameters of the universe in the redshift. The result is
@@
In deriving relation (111c), it is utilised that, subsequent to decoupling, $T_{{\scriptscriptstyle(0)}\gamma}\propto a^{-1}$. The dimensionless time, defined as $\tau:=t/t_{\text{dec}}$, is given by
@@
where relation (111b) with $z(t_{\text{p}})=0$ is used. Upon substituting the result (92) into expression (107) and utilizing the relations (111), the following is obtained:
where it is assumed that at decoupling, the matter and radiation temperatures were equal, that is, $T_{\scriptscriptstyle(0)}(t_{\text{dec}})=T_{{\scriptscriptstyle(0)}\gamma}(t_{\text{dec}})$. In this manner, the parameter $\mu_{\text{m}}$ has been expressed in observable quantities. Upon substituting the numerical values from Tables 1 and 2, one obtains the result:
It can thus be concluded that the parameter $\mu_{\text{m}}$ is dependent solely on the initial scale $\lambda_{\text{dec}}$ of a density fluctuation.
Table 1: Planck satellite results
$z(t_{\text{dec}})$ = 1090
$z(t_{\text{eq}})$ = 3387
$cH(t_{\text{p}})$ = $67.66\,\text{km}\,\text{s}^{-1}\text{Mpc}^{-1}$
[…]
$\lvert\delta_{T_{\gamma}}(t_{\text{dec}},\bm{x})\rvert$ $\lesssim$ $10^{-5}$
XII.2.2 Initial Values
In order to solve Eq. (108), it is necessary to determine the initial values of the quantities $\delta_{\varepsilon}$ and $\delta^{\prime}_{\varepsilon}$. Furthermore, it is imperative to ascertain the values of the local nonadiabatic random pressure fluctuation $\delta_{T}$ in expression (103).
Planck Satellite
The Planck observations [43] of the fluctuations $\delta_{T_{\gamma}}(t_{\text{dec}},\bm{x})$ in the background radiation temperature imply that $\lvert\delta_{\varepsilon}(t_{\text{dec}},\bm{x})\rvert\lesssim 10^{-5}$. In the absence of knowledge regarding the initial growth rate, it is assumed that $\dot{\delta}_{\varepsilon}(t_{\text{dec}},\bm{x})\approx 0$. Therefore, the initial values for Eq. (108) are as follows:
As demonstrated in the calculations presented in Sec. XII.2.3, the outcome of Eq. (108) is largely independent of the initial value for $\delta_{\varepsilon}$, provided that it satisfies the condition of having a maximum value of $\lvert\delta_{\varepsilon}(t_{\text{dec}},\bm{q})\rvert\leq 10^{-4}$.
Nonadiabatic Pressure Fluctuations
The values of the quantities $\delta_{\varepsilon}(t,\bm{q})$ and $\delta_{n}(t,\bm{q})$ are related to the local nonadiabatic random pressure fluctuation $\delta_{T}(t_{\text{dec}},\bm{q})$ by Eq. (101). Immediately following the decoupling, this equation reads:
$\delta_{n}(t_{\text{dec}},\bm{q})-\delta_{\varepsilon}(t_{\text{dec}},\bm{q})\approx-4.1\times 10^{-10}\delta_{T}(t_{\text{dec}},\bm{q}),$ (116)
where it has been assumed that at decoupling, the matter temperature was equal to the radiation temperature, $T_{\scriptscriptstyle(0)}(t_{\text{dec}})=T_{{\scriptscriptstyle(0)}\gamma}(t_{\text{dec}})$. Expressions (91) and (111c) were employed in conjunction with the values presented in Tables 1 and 2.
The transition of the universe from the era preceding decoupling to the era following decoupling was rapid and chaotic. A notable decline was observed in the mean particle velocity, and the pressure. Furthermore, the values of both $\delta_{\varepsilon}(t_{\text{dec}},\bm{q})$ and $\delta_{n}(t_{\text{dec}},\bm{q})$ exhibited minor irregularities on the surface of last scattering. Therefore, it is probable that after the transition, the initial values of the difference between fluctuations, represented by ${(\delta_{n}-\delta_{\varepsilon})(t_{\text{dec}},\bm{q})}$, were randomly distributed among all density fluctuations. From expression (116), it can be seen that very small differences between the initial energy density and particle number density fluctuations result in relatively large positive or negative fluctuations in the nonadiabatic component of the pressure fluctuations in expression (103). Therefore, the nonadiabatic pressure fluctuations, represented by $\delta_{T}(t_{\text{dec}},\bm{q})$, are randomly (102) distributed among the multitude of density fluctuations.
@@ XII.2.2 Initial Values
XII.2.3 Structure Formation

### v3

@@ XIII.2 Era after Decoupling of Matter and Radiation
In these expressions, the symbol $k_{\text{B}}$ represents Boltzmann’s constant, the quantity $m$ denotes the proton rest mass, and the temperature of the matter is indicated by the symbol $T$. The rest mass energy density is $nmc^{2}$ and the kinetic energy density is $\tfrac{3}{2}nk_{\text{B}}T$.
The elimination of the temperature $T$ from the equations of state (89), yields the equation of state for the pressure [see Ref. [42], Eq. (2.10.27)]:
@@
The partial derivatives are determined using the expressions (14b) and are given by
@@
$\quad\beta\approx\dfrac{v_{\text{s}}}{c}=\sqrt{\dfrac{5}{3}\dfrac{k_{\text{B}}T_{\scriptscriptstyle(0)}}{mc^{2}}},$ (93)
where $v_{\text{s}}$ represents the adiabatic speed of sound.
From $\beta^{2}\approx\tfrac{5}{3}w$, we find from Eq. (44) that $\dot{w}\approx-2Hw$. Therefore, with $H:=\dot{a}/a$, it can be concluded that $w\propto a^{-2}$. Given that $w$ is proportional to $T_{{\scriptscriptstyle(0)}}$, the well-known result (see Ref. [42], Eq. (15.5.16) with $\gamma=\tfrac{5}{3}$, where $\gamma:=c_{\text{p}}/c_{\text{v}}$} is the adiabatic index) is obtained:
@@
The subsequent step is to derive the evolution equations for relative density perturbations. The proportionality (94) implies with (93) that $\dot{\beta}/\beta=-H$. The system of equations (62) with coefficients (76) can now be rewritten in the following form:
@@
where it is used that $w$ and $\beta^{2}$ are negligible with respect to constants of order one.
@@
$(\delta_{n}-\delta_{\varepsilon})\propto a^{-2}.$ (96)
In order to relate this proportionality to thermodynamic quantities, it is necessary to express the perturbed equation of state for the energy density in terms of relative density perturbations. It can be deduced from $\varepsilon=\varepsilon(n,T)$ that
@@
$\varepsilon_{{\scriptscriptstyle(1)}}^{{\textnormal{phys}}}=\left(\dfrac{\partial\varepsilon}{\partial n}\right)_{T}n_{\scriptscriptstyle(1)}^{{\textnormal{phys}}}+\left(\dfrac{\partial\varepsilon}{\partial T}\right)_{n}T_{\scriptscriptstyle(1)}^{{\textnormal{phys}}},$ (98)
where expressions (6a) and (72) have been used. By using the expression for $\varepsilon$ in (89) to eliminate the partial derivatives, one obtains the perturbed equation of state:
@@
Upon dividing (99) by $\varepsilon_{{\scriptscriptstyle(0)}}=n_{{\scriptscriptstyle(0)}}mc^{2}+\tfrac{3}{2}n_{{\scriptscriptstyle(0)}}k_{\text{B}}T_{{\scriptscriptstyle(0)}}$ and using the exact value of the ratio $w:=p_{{\scriptscriptstyle(0)}}/\varepsilon_{{\scriptscriptstyle(0)}}$ (92), the perturbed equation of state for the energy density expressed in relative density perturbations (61) is obtained:
@@
The quantity $\delta_{T}$ is defined by $\delta_{T}:=T^{\textnormal{phys}}_{\scriptscriptstyle(1)}/T_{\scriptscriptstyle(0)}$. The solution to Eq. (95b) is obtained by combining (96) and (100). Using the approximate value of $w$ implies that $w\propto a^{-2}$ [see (94)], making $\delta_{T}$ independent of time. Therefore, the solution to Eq. (95b) is given by
@@
The quantity $\delta_{T}(t_{\text{dec}},\bm{x})$ represents the relative temperature perturbation of matter at the time $t_{\text{dec}}$, which marks the decoupling of matter from radiation.
According to (71), (92) and (101) the entropy per particle is:
@@
Thus, the quantity $\delta_{T}$ is random. As a result, it is generally not equal to zero, which implies that density perturbations are generally nonadiabatic.
Using (91), (93), and (101), the local relative pressure perturbation $\delta_{p}:=p^{\textnormal{phys}}_{\scriptscriptstyle(1)}/p_{\scriptscriptstyle(0)}$ can be calculated from expression (74). One obtains
@@
In this context, the terms $\tfrac{5}{3}\delta_{\varepsilon}(t,\bm{x})$ and $\delta_{T}(t_{\text{dec}},\bm{x})$ represent the adiabatic and the random, nonadiabatic relative pressure perturbation, respectively.
The solution (101) to Eq. (95b) allows for the reformulation of the second-order equation (95a). To that end, the solutions to the background equations (75) are required. Given that the pressure with respect to the rest mass energy density can be disregarded ($w\ll 1$), the following solutions are obtained:
@@
The dimensionless time, denoted by the symbol $\tau$, is defined by $\tau:=t/t_{\text{dec}}\geq 1$. This definition implies that
@@
By using Eqs. (75a) and (101), as well as the expressions (93), (104), (105), the Helmholtz equation $\nabla^{2}\delta=-|\bm{q}|^{2}\delta$, and $w\propto a^{-2}$, it can be demonstrated that Eq. (95a) can be expressed in the following manner:
$\delta_{\varepsilon}^{\prime\prime}+\dfrac{2}{\tau}\delta_{\varepsilon}^{\prime}+\left(\dfrac{4}{9}\dfrac{\mu_{\text{m}}^{2}}{\tau^{8/3}}-\dfrac{10}{9\tau^{2}}\right)\delta_{\varepsilon}=-\dfrac{4}{15}\dfrac{\mu^{2}_{\text{m}}}{\tau^{8/3}}\delta_{T}(t_{\text{dec}},\bm{q}),$ (106)
where the prime denotes differentiation with respect to $\tau$. With $\lvert\bm{q}\rvert=2\pi/\lambda_{\text{dec}}$, the parameter $\mu_{\text{m}}$ is defined by
@@
In this expression, $\lambda_{\text{dec}}:=\lambda a(t_{\text{dec}})$ represents the physical scale of a perturbation at $t=t_{\text{dec}}$, and $v_{\text{s}}(t_{\text{dec}})/c\approx\beta(t_{\text{dec}})$ is given by (93).
Finally, the expression (103) enables the reformulation of Eq. (106) in a form suitable for the study of the evolution of relative density perturbations:
@@
For large-scale perturbations, i.e., when $\lambda_{\text{dec}}\rightarrow\infty$, implying that $\mu_{\text{m}}\rightarrow 0$, relative pressure perturbations $\delta_{p}$ played a negligible role in their evolution. Therefore, the solution to Eq. (108) is
@@
The solution proportional to $\tau^{-5/3}$ dies out quickly. Thus, large-scale relative density perturbations evolve nearly proportionally to $\tau^{2/3}$, which is close to the solution (60) for a pressureless fluid. However, since $\vartheta_{\scriptscriptstyle(1)}^{\textnormal{phys}}(\tau)=-\tfrac{7}{9}\tau^{-4/3}$, (158b), there is a pressure-gradient-driven fluid flow. Therefore, the solution (109) is invalid for a pressureless fluid. As shown in Sec. X, the equations for scalar perturbations (47) are, for a pressureless fluid, reduced to a single equation, namely the time-independent Poisson equation (56).
A comparison of the solutions (109) and (158a) reveals that the solution proportional to $\tau^{2/3}$ in (109) is also a solution to the homogeneous part of Eq. (157a). The solution proportional to $\tau^{-5/3}$ is the particular solution to the inhomogeneous equation. This solution can only be obtained if the covariant divergence $\vartheta_{{\scriptscriptstyle(1)}}$ is taken into account, as demonstrated in Appendix C.2. Since in the standard equation (156) $\vartheta_{{\scriptscriptstyle(1)}}=0$ (i.e., no fluid flow), it does not produce growing density perturbations.
The presence of the gauge constant $C$, precludes the derivation of solution (109) from the system (157).
XIII.2.1 Cosmological Quantities
Table 1: Planck satellite resultsy
$z(t_{\text{dec}})$ = 1090
$z(t_{\text{eq}})$ = 3387
$cH(t_{\text{p}})$ = $67.66\,\text{km}\,\text{s}^{-1}\text{Mpc}^{-1}$
$T_{{\scriptscriptstyle(0)}\gamma}(t_{\text{p}})$ = $2.725\,\text{K}$
$t_{\text{p}}$ = $13.79\,\text{Gyr}$
$\lvert\delta_{T_{\gamma}}(t_{\text{dec}},\bm{x})\rvert \lesssim$ $10^{-5}$
To investigate the evolution of relative density perturbations using Eq. (108), the parameter $\mu_{\text{m}}$ (107) must first be expressed in terms of observable quantities, namely the redshift $z(t_{\text{dec}})$ at the time of decoupling, the present value of the Hubble parameter $cH(t_{\text{p}})$, and the present value of the temperature $T_{{\scriptscriptstyle(0)}}(t_{\text{p}})$ of the cosmological fluid.
@@ XIII.2.1 Cosmological Quantities
where $a(t_{\text{p}})$ is the present value of the scale factor. The Friedmann equation (11a) shows that, for a flat flrw universe, one may set $a(t_{\text{p}})=1$. Using the background solutions (104), it is possible to express the relevant parameters of the universe in the redshift:
@@
In deriving relation (111c), it is used that, subsequent to decoupling, $T_{{\scriptscriptstyle(0)}\gamma}\propto a^{-1}$. Upon substituting (93) into expression (107) and using the relations (111), we obtain
where it is assumed that at decoupling, the matter and radiation temperatures were equal, that is, $T_{\scriptscriptstyle(0)}(t_{\text{dec}})=T_{{\scriptscriptstyle(0)}\gamma}(t_{\text{dec}})$. In this manner, the parameter $\mu_{\text{m}}$ has been expressed in observable quantities. Upon substituting the numerical values from Tables 1 and 2, one obtains
It can thus be concluded that the parameter $\mu_{\text{m}}$ is dependent solely on the initial scale $\lambda_{\text{dec}}$ of a density perturbation.
Finally, the dimensionless time, defined as $\tau:=t/t_{\text{dec}}$, is expressed in terms of the redshift, given by:
@@
where relation (111b) with $z(t_{\text{p}})=0$ is used.
@@
XIII.2.2 Initial Values
In order to solve Eq. (108), it is necessary to determine the initial values of the quantities $\delta_{\varepsilon}$ and $\delta^{\prime}_{\varepsilon}$. In addition, values of the local random, nonadiabatic relative pressure perturbation $\delta_{T}$ in expression (103), are required.
Planck Satellite
The Planck observations [50] of the relative perturbations $\delta_{T_{\gamma}}(t_{\text{dec}},\bm{x})$ in the background radiation temperature imply that $\lvert\delta_{\varepsilon}(t_{\text{dec}},\bm{x})\rvert\lesssim 10^{-5}$. In the absence of knowledge regarding the initial growth rate, it is assumed that $\dot{\delta}_{\varepsilon}(t_{\text{dec}},\bm{x})\approx 0$. Therefore, the initial values for Eq. (108) are as follows:
As demonstrated in the calculations presented in Sec. XIII.2.3, the outcome of Eq. (108) is almost entirely independent of the initial value for $\delta_{\varepsilon}$, provided that it satisfies the condition $\lvert\delta_{\varepsilon}(t_{\text{dec}},\bm{q})\rvert\leq 10^{-4}$.
Random Nonadiabatic Pressure Perturbations
The initial values $\delta_{\varepsilon}(t_{\text{dec}},\bm{q})$ and $\delta_{n}(t_{\text{dec}},\bm{q})$ are related to the local random, nonadiabatic relative pressure perturbation $\delta_{T}(t_{\text{dec}},\bm{q})$ by expression (101). Since $w\ll 1$}, very small values of the difference $\delta_{n}(t_{\text{dec}},\bm{q})-\delta_{\varepsilon}(t_{\text{dec}},\bm{q})$ result in relatively large positive or negative values of $\delta_{T}(t_{\text{dec}},\bm{q})$.
The fact that $\delta_{T}(t_{\text{dec}},\bm{q})$ takes on random values follows from expression (102). However, using Eq. (101) the randomness of $\delta_{T}(t_{\text{dec}},\bm{q})$ can also be explained as follows. The transition of the universe from the era preceding decoupling to the era following decoupling was rapid and chaotic. A notable decline was observed in the mean particle velocity, and the pressure. As a consequence, the values of both $\delta_{\varepsilon}(t_{\text{dec}},\bm{q})$ and $\delta_{n}(t_{\text{dec}},\bm{q})$ exhibited minor irregularities on the surface of last scattering. Therefore, it is probable that after the transition, the initial values of the difference between the relative density perturbations, represented by $(\delta_{n}-\delta_{\varepsilon})(t_{\text{dec}},\bm{q})$}, were randomly distributed among all density perturbations.
XIII.2.3 Structure Formation